Given that ​f(x)=4x+1 and ​g(x)=x^2-3x-4​, find ​(g*f)(6​)
1
Expert's answer
2019-10-07T09:35:40-0400

g*f=(x^2 - 3x - 4)(4x+1) = 4x3 - 11x2 - 19x - 4

g*f(6) = 4 * 63 - 11 * 62 - 19 * 6 - 4 = 350